# breaklines Repository-Level Compositional Code Translation and Validation

### v2

@@ 7.3.2. Translation Bugs
Our publicly available artifacts (Lab, 2024) contain partial translations and fixed versions as separate commits. These commits can serve as useful benchmarks for evaluating techniques such as fault localization, program repair, and test generation. This section shows four instances of such translation bugs. The two most prevalent sources of translation bugs are mismatches between APIs and behavioral differences in the PLs. The code snippet below demonstrates a bug that happened due to a mismatch in the logic of Calendar (Java) and datetime (Python). Line in Java sets the MONTH field to 0, which corresponds to the first month of the year (January). Similarly, the Python translation sets the month attribute to 0; however, in the Python library, January is the first month, i.e., the correct translation should use index 1.
@@
The last code snippet shows an example where the Java behavior of an iterator is unavailable in Python. Specifically, the incorrect Python translation uses next() to implement both .next() and .hasNext() calls of Java iterator. The issue is that calling next() increments the iterator in Python. The correct translation should implement PeekableIterator interface in Python with a method hasNext() -> bool.
@@ 7.4. RQ3: Impact of Test Decomposition
We previously showed the effectiveness of test translation in validating the runtime behavior or even the functional correctness of application method fragments (§7.2.2). To better understand the effect of test decomposition and how it helps with the test translation coupling effect, we collected translated unit tests with at least two corresponding decomposed test fragments. We further filtered out these unit tests and kept those that all their decomposed test fragments were executed, regardless of passing or failing. The yellow bars in Figure 6 show the percentage of our selected unit tests from the translated test suites. For Commons-CSV and Commons-Validator, none of the unit tests met the criteria, so we excluded them from further investigation.
We categorized the selected unit tests into two groups: Those with all their test fragments pass (green bars in Figure 6), and those with at least one test fragment fail (red bars in Figure 6). For the unit tests in the latter group, we calculate the pass rate of the decomposed test fragments. The blue box chart in Figure 6 shows the distribution of the measured pass rate per unit test. These unit tests would have been marked as fail without test decomposition. However, we can observe that these tests can be decomposed into test fragments that of them pass, on average. These results confirm how decomposed test fragments were useful in helping developers localize translation bugs more easily and resolve the translation bugs faster.
@@ 7.5. RQ4: Impact of Test Coverage
While existing developer-written tests are useful for validating functional equivalence, they can pose two major issues for automated code translation and validation. First, the coverage for these tests can be extremely low (e.g., for Commons-Pool (Foundation, 2024g)), preventing most of the code from being validated (as shown in RQ1, the translation validation rate strongly correlates with test suites’ (method) coverage). Second, executing a developer-written test can have a long call sequence. To show the positive impact of better, more focused tests with higher coverage on translation validation, we automatically generated more tests using EvoSuite (Fraser and Arcuri, 2011). We generated EvoSuite tests with the test generator’s default settings: used DynaMOSA (Panichella et al., 2018) as optimization algorithm and set a timeout of seconds as stopping criteria.
Table 3 compares the properties of the developer-written tests and EvoSuite tests. Since we configure EvoSuite to generate several tests per each Java class, the average Method Coverage of its tests is higher than developer-written tests ( compared to ). Furthermore, the average number of methods executed per single test is almost half that of decomposed test suites ( compared to methods). We translated EvoSuite tests and executed the translated projects using them. As corroborated by the numbers under columns TPR, we can see that we achieve a higher test pass rate ( compared to ) on the same translated code. Note that not all the EvoSuite tests have assertions, and even if they do, the quality of the assertions could be lower compared to developer-written tests (e.g., checking trivial properties that result in a test pass). Thereby, we only claim that higher TPR of such tests enhance runtime behavior validation, which is still promising in code translation. Unfortunately, EvoSuite is not compatible with Java , which is required for the GraalVM component of AlphaTrans. Otherwise, we could use the generated tests in AlphaTrans. We anticipate including it in the loops could have also improved the overall quality of translations.
| Subjects | Developer-Written Test | Evosuite Test | |||||||||||||||||||
@@
| cli | 93.4 | 3036 | 34.3 | 10.1 | 83.5 | 569 | 12.2 | 13.9 | |||||||||||||
| codec | 86.5 | 3522 | 10.6 | 9.4 | 86.9 | 1141 | 8.0 | 48.6 | |||||||||||||
| csv | 87.7 | 1219 | 52.6 | 0.2 | 68.5 | 220 | 39.2 | 8.6 | |||||||||||||
| exec | 25.4 | 311 | 19.0 | 19.3 | 76.8 | 245 | 6.3 | 28.2 | |||||||||||||
| fast-pfor | 61.9 | 249 | 41.6 | 20.1 | 82.6 | 1843 | 4.3 | 30.9 | |||||||||||||
| fileupload | 50.5 | 93 | 3.5 | 63.4 | 84.4 | 231 | 5.3 | 41.1 | |||||||||||||
| graph | 58.4 | 933 | 25.0 | 11.0 | 86.3 | 800 | 9.0 | 9.1 | |||||||||||||
| jansi | 24.6 | 187 | 13.6 | 1.1 | 70.3 | 332 | 9.1 | 14.2 | |||||||||||||
| pool | 20.3 | 287 | 6.5 | 6.6 | 69.1 | 394 | 7.4 | 5.6 | |||||||||||||
| validator | 53.2 | 1479 | 11.7 | 11.7 | 84.5 | 1305 | 13.4 | 3.3 | |||||||||||||
56.2 | 11316 | 21.8 | 15.3 | 79.3 | 7080 | 11.4 | 20.3 | |||||||||||||
@@ 8. Related Work
Code translation using non-LLM-based approaches
@@
Code translation using LLMs
@@ 9. Threats to Validity
External Validity. One of the key external threats is the generalizability of our approach. The translation pipeline is very generic and can be extended for more language pairs. Also, the majority of the tools that we used support a large set of programming languages. However, to expedite the research in repository-level translation, we build the first version of AlphaTrans to be specific towards Java to Python.
Internal Validity. One major internal threat can be that the results are calculated based on a single run. To mitigate that threat, we ran AlphaTrans with greedy decoding and temperature , which reduces the model’s creativity but makes the output consistent with several runs. Another threat in this group can be the manual validation of the translated types. To address that, several authors have verified the types individually and consulted API documents when necessary.
@@ 10. Concluding Remarks
In this paper, we introduced AlphaTrans, a neuro-symbolic approach that combines the power of static analysis and emerging abilities of Code LLMs in code synthesis to automate repository-level code translation and validation. AlphaTrans decomposes the program into smaller fragments and translates the fragments in the reverse call order, originally building the project in the target language. In addition to syntax check, AlphaTrans implements two levels of validation through GraalVM and test translation. Our results demonstrate the effectiveness of AlphaTrans in translating ten real-world Java projects to Python, achieving syntactical correctness, runtime behavior validation, and functional equivalence. AlphaTrans is the first approach to translate and validate the entire repository, and we envision several research directions to advance repository-level code translation and validation as follows:
One of the major challenges of repository-level code translation is identifying suitable library APIs in the target PL. Often, equivalent Python APIs may not exist, requiring new code generation or translation of the library API itself. Even if similar libraries exist, the logic of libraries might be different in two PLs. AlphaTrans supports translating frequently used APIs and aims to build a generic pipeline. Supporting all the libraries in the pipeline remains an open challenge. Furthermore, while the idea of compositional translation and validation is PL-agnostic, the static analysis makes the extension of AlphaTrans to translating from other source projects challenging. Devising LLM-enabled or PL-agnostic static analysis approaches can benefit code translation approaches such as AlphaTrans. We also showed that the quality of the source project test suite can significantly impact the translation validation results. As part of our future work, we plan to integrate an LLM-based test generator into the AlphaTrans pipeline to advance the validation component.

### v3

@@ 7.3.2. Translation Bugs
Our artifacts (Lab, 2025) contain partial translations and fixed versions as separate commits. These commits can serve as useful benchmarks for evaluating fault localization, program repair, and test generation techniques. This section shows four instances of such translation bugs.
The two most prevalent sources of translation bugs are mismatches between APIs and behavioral differences in the PLs. The code snippet below demonstrates a bug that happened due to a mismatch in the logic of Calendar (Java) and datetime (Python). Line in Java sets the MONTH field to 0, which corresponds to the first month of the year (January). Similarly, the Python translation sets the month attribute to 0; however, in the Python library, January is the first month, i.e., the correct translation should use index 1.
@@
The last code snippet demonstrates the Java behavior of an iterator that is unavailable in Python. The incorrect Python translation uses next() to implement both .next() and .hasNext() APIs. However, calling next() increments the iterator in Python. The correct translation should implement PeekableIterator interface in Python with a method hasNext() -> bool.
@@ 7.4. RQ3: Impact of Test Decomposition
We previously showed the effectiveness of test translation in validating the runtime behavior or functional correctness of application method fragments (§7.2.2). To better understand how test decomposition helps test translation coupling effect, we collected translated unit tests with at least two corresponding decomposed test fragments. We further filtered out these unit tests and kept those that all their decomposed test fragments were executed, regardless of passing or failing. The yellow bars in Figure 6 show the percentage of our selected unit tests from the translated test suites. None of the tests met the criteria for Commons-CSV, so we excluded it from further investigation.
We categorized the selected unit tests into two groups: Those with all their test fragments pass (green bars in Figure 6), and those with at least one test fragment fail (red bars in Figure 6). For the unit tests in the latter group, we calculate the pass rate of the decomposed test fragments. The blue box chart in Figure 6 shows the distribution of the measured pass rate per unit test. These unit tests would have been marked as fail without test decomposition. However, we can observe that these tests can be decomposed into test fragments that of them pass. These results confirm how decomposed test fragments were useful in helping developers localize translation bugs more easily and resolve the translation bugs faster.
@@ 7.5. RQ4: Impact of Test Coverage
While existing developer-written tests are useful for validating functional equivalence, they can pose two major issues for automated code translation and validation. First, the coverage for these tests can be extremely low (e.g., for Commons-FileUpload (Foundation, 2024e) reported in Table 1), preventing most of the code from being validated (as shown in RQ1, the translation validation rate strongly correlates with test suites’ (method) coverage). Second, executing a developer-written test can have a long call sequence. To show the positive impact of better, more focused tests with higher coverage on translation validation, we automatically generated additional tests using EvoSuite (Fraser and Arcuri, 2011). We used DynaMOSA (Panichella et al., 2018) as an optimization algorithm and set a timeout of seconds as a stopping criterion for the generation of evolutionary tests.
| Developer-Written Test | EvoSuite Test | |||||||||||||||||||||
| Subjects |
ATP (%) |
@@
TPR+ (%) | ATP+ (%) | ||||||||||||
| cli | 94.14 | 3036 | 34.25 | 10.08 | 8.42 | 95.97 | 569 | 12.15 | 2.99 | 1.47 | ||||||||||||
| codec | 91.03 | 3522 | 10.56 | 9.43 | 4.12 | 80.74 | 1141 | 8.02 | 3.51 | 0.88 | ||||||||||||
| csv | 90.64 | 1219 | 52.62 | 0 | 0 | 74.04 | 220 | 39.16 | 0 | 0.00 | ||||||||||||
| exec | 54.84 | 311 | 18.99 | 19.29 | 4.44 | 61.29 | 245 | 6.32 | 3.27 | 1.21 | ||||||||||||
| fast-pfor | 54.55 | 249 | 41.62 | 20.08 | 4.28 | 39.17 | 1843 | 4.31 | 5.59 | 1.07 | ||||||||||||
| fileupload | 13.02 | 93 | 3.54 | 63.44 | 3.65 | 70.31 | 231 | 5.29 | 11.26 | 2.60 | ||||||||||||
| graph | 58.78 | 933 | 25.02 | 11.04 | 0.00 | 76.71 | 800 | 9.00 | 5.13 | 0.92 | ||||||||||||
| jansi | 23.47 | 187 | 13.57 | 1.07 | 0.24 | 51.83 | 332 | 9.08 | 3.31 | 0.73 | ||||||||||||
| pool | 22.29 | 287 | 6.52 | 6.62 | 1.61 | 37.24 | 394 | 7.36 | 10.41 | 2.20 | ||||||||||||
| validator | 63.31 | 1479 | 11.68 | 11.70 | 3.25 | 81.42 | 1305 | 13.43 | 9.73 | 7.59 | ||||||||||||
| Total | 56.57 | 11316 | 21.84 | 9.76 | 2.88 | 66.87 | 7080 | 11.41 | 5.85 | 2.11 | ||||||||||||
Table 3 compares the developer-written and EvoSuite tests’ properties. Evosuite tests cover more methods than the developer-written test suite ( Method Coverage compared to ). Furthermore, the average number of methods executed per single test is almost half that of decomposed test suites ( compared to methods). To demonstrate the impact of test quality on AlphaTrans’s overall performance, we translated and executed the EvoSuite-generated tests. Corroborated by the numbers under TPR+ and ATP+ columns, we can see that augmenting the test suite increases the method coverage, and thereby, the TPR and ATP numbers from RQ1 by and , respectively. Not all the EvoSuite tests have assertions, and even if they do, the quality of the assertions could be lower compared to developer-written tests (e.g., checking trivial properties that result in a test pass). Thereby, we only claim that higher TPR and ATP of such tests enhance runtime validation, which is still promising in code translation. EvoSuite is incompatible with Java , and hence GraalVM, preventing us from using Evosuite-generated tests in AlphaTrans. We anticipate that including it in the loops could have also improved the overall quality of translations.
7.6. RQ5: Ablation Study
| GraalVM | Test Translation | |||||||||||
| Subjects | AMF | Syntax Check (%) | SNEF (%) | GS (%) | GF (%) | GE (%) | TNEF (%) | ATP (%) | OTF (%) | MTF (%) | ATF (%) | TPR (%) |
| cli | 276 | 100 | 5.80 | 0 | 0 | 94.20 | 85.51 | 0 | 2.17 | 1.09 | 5.43 | 0.46 |
| codec | 678 | 98.82 | 10.62 | 0 | 0 | 89.38 | 72.86 | 2.06 | 5.46 | 2.80 | 6.19 | 5.04 |
| csv | 235 | 100 | 8.51 | 0 | 0 | 91.49 | 88.09 | 0.43 | 0.43 | 0.85 | 1.70 | 0.32 |
| exec | 253 | 100 | 46.25 | 26.88 | 4.35 | 22.53 | 49.41 | 1.58 | 1.19 | 0 | 1.58 | 4.29 |
| fast-pfor | 754 | 98.14 | 46.02 | 0 | 0 | 53.98 | 38.20 | 0.93 | 6.63 | 0.80 | 7.43 | 4.88 |
| fileupload | 168 | 100 | 85.12 | 10.71 | 0.60 | 3.57 | 8.93 | 1.19 | 2.98 | 1.19 | 0.60 | 23.08 |
| graph | 556 | 99.82 | 40.65 | 0 | 0 | 59.35 | 56.12 | 0 | 1.98 | 0.36 | 0.90 | 5.48 |
| jansi | 314 | 99.68 | 69.75 | 0 | 0 | 30.25 | 29.94 | 0 | 0.32 | 0 | 0 | 0 |
| pool | 574 | 99.83 | 70.56 | 0 | 0 | 29.44 | 27.00 | 1.39 | 0 | 0.35 | 0.70 | 5.48 |
| validator | 623 | 99.20 | 33.39 | 18.78 | 20.06 | 27.77 | 66.29 | 0 | 0 | 0.32 | 0 | 0.43 |
| Total | 4431 | 99.32 | 40.01 | 4.58 | 3.09 | 52.31 | 52.79 | 0.81 | 2.57 | 0.86 | 2.96 | 3.05 |
We performed three ablation studies to investigate the impact of program transformation, choice of LLM, and program decomposition on the performance of AlphaTrans.
7.6.1. Impact of Program Transformation
We removed the program transformation component of AlphaTrans and executed the entire pipeline. The results in Table 4 show that without transformation (e.g., resolving method/constructor overloading), the performance of AlphaTrans drops significantly: GS, ATP, and TPR values decreased to (from ), (from ), and (from ), respectively. This is because Python does not support overloading and only considers the last method/constructor implementation, resulting in runtime errors or test failures. Also, GE values increase due to the interference of overloaded code constructs with GraalVM.
| GraalVM | Test Translation | |||||||||||||||||||
| OTF (%) | MTF (%) | ATF (%) | M1 | |||||||||||||||||
| Subjects | Syntax Check (%) | SNEF (%) | GS (%) | GF (%) | GE (%) | TNEF (%) | ATP (%) | O | RE | AF | O | RE | AF | O | RE | AF | TPR (%) | All | Some | Cost ($) |
| cli | 99.63 | 5.86 | 76.92 | 8.79 | 8.42 | 78.39 | 4.76 | 3.66 | 90.00 | 10.00 | 6.23 | 100 | 0 | 1.10 | 100 | 0 | 2.47 | 0 | 1 | 19.69 |
| codec | 97.79 | 8.97 | 42.50 | 28.53 | 20.00 | 74.56 | 3.82 | 4.12 | 78.57 | 21.43 | 6.18 | 30.62 | 69.38 | 2.35 | 73.33 | 26.67 | 9.57 | 2 | 18 | 39.97 |
| csv | 98.30 | 9.36 | 41.28 | 25.96 | 23.40 | 86.81 | 0 | 0 | 0 | 0 | 1.70 | 78.17 | 21.83 | 2.13 | 96.72 | 3.28 | 0.98 | 0 | 8 | 16.66 |
| exec | 99.60 | 45.16 | 35.89 | 2.42 | 16.53 | 44.76 | 4.84 | 0 | 0 | 0 | 5.24 | 0 | 100 | 0 | 0 | 0 | 28.62 | 1 | 1 | 3.90 |
| fast-pfor | 97.46 | 45.45 | 14.71 | 16.98 | 22.86 | 52.94 | 1.47 | 0 | 0 | 0 | 0 | 0 | 0 | 0.13 | 100 | 0 | 2.41 | 2 | 0 | 12.93 |
| fileupload | 99.48 | 86.98 | 9.90 | 0.52 | 2.60 | 4.17 | 6.77 | 1.56 | 0 | 100 | 0.52 | 100 | 0 | 0 | 0 | 0 | 54.84 | 1 | 3 | 2.25 |
| graph | 98.71 | 41.22 | 27.73 | 19.04 | 12.01 | 58.78 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 12.03 |
| jansi | 99.76 | 76.53 | 8.07 | 11.98 | 3.42 | 23.47 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 3.66 |
| pool | 99.71 | 77.71 | 7.48 | 1.32 | 13.49 | 21.70 | 0.59 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2.09 | 0 | 0 | 7.33 |
| validator | 99.23 | 36.69 | 38.24 | 15.94 | 9.13 | 54.95 | 4.02 | 2.32 | 93.33 | 6.67 | 1.55 | 70.59 | 29.41 | 0.46 | 100 | 0 | 10.41 | 0 | 3 | 25.53 |
| Total | 98.80 | 43.43 | 27.83 | 14.54 | 14.20 | 50.64 | 2.26 | 1.20 | 80.36 | 19.64 | 1.87 | 59.39 | 40.61 | 0.60 | 86.59 | 13.41 | 6.45 | 6 | 34 | 143.95 |
7.6.2. Choice of LLM
For this experiment, we replaced the DeepSeek-Coder-33b-Instruct with GPT-4o and repeated the entire pipeline of AlphaTrans (Table 5). A stronger model such as GPT-4o improves the translation quality —functional equivalence increases from to . For some projects, the ATP and TPR rates are higher for DeepSeek-Coder-33b-Instruct translations, while for the others, GPT-4o results in higher values. We investigated each LLM’s successful AMF translations to better understand the differences. We observed a huge overlap between successful AMFs and the unique benefits each LLM provides in code translation (Figure 7). GPT-4o handles API translation and type casting better, resolving the first three translation bugs discussed in §7.3. In contrast, it tends to add unnecessary code, mostly due to error handling, which results in a functional mismatch. In the example below, create2 method can take a None variable, and its implementation performs error handling when necessary. GPT-4o adds unnecessary error handling code, interfering with program logic and resulting in test failures.
[frame=lines,framesep=1mm,baselinestretch=0.5, fontsize=, breaklines, breakanywhere, linenos,numbersep=2pt]java ———– JAVA SOURCE CODE ———– public Option create1(OptionGroup optGroup) return create2(optGroup);
[escapeinside=——, frame=lines,framesep=1mm,baselinestretch=0.5, fontsize=, breaklines, breakanywhere, linenos,numbersep=2pt, highlightlines=3,4]py ———- PYTHON TRANSLATION ———- def create1(self, optGroup: OptionGroup) -¿ Option: ——+ if optGroup is None: ——+ raise ValueError(”optGroup is None”) return self.create2(optGroup)
It is worth noting that using commercial LLMs comes at a cost. The last column of Table 5 (column Cost) shows the cost of using GPT-4o for repeating the experiments, resulting in the total cost of for translating all the subjects (average cost of per project).
7.6.3. Impact of Program Decomposition
For this ablation study, we prompted GPT-4o and DeepSeek-Coder-33b-Instruct file-by-file and evaluated translation correctness through test execution and GraalVM (Table 6). Not surprisingly, a considerable percentage of the files exceeded the model context window size, particularly for DeepSeekCoder, with of the files encountering this problem. Among the prompted files, ( for DeepSeekCoder and for GPT-4o) were syntactically incorrect. Translations that passed the syntactic correctness check did not pass any translated test execution, whereas Graal validated of the files in total ( for DeepSeekCoder and for GPT-4o). Note that file-level Graal validation does not mean that all the methods are correctly translated and validated; only the methods in the class are executed by tests. Manual analysis of the results of this experiment revealed that one prominent culprit of test failures was LLM hallucination with method/variable names. Given that we had no skeleton construction in this baseline, such issues could not be avoided; this demonstrates the usefulness of skeleton construction and incremental translation in AlphaTrans.
@@ 8. Related Work
8.1. Code translation using non-LLM-based approaches.
@@
| GPT-4o | DeepSeek-Coder | ||||||||
| Subjects | # Files | Over Context | Syntax Error | GS | TPR (%) | Over Context | Syntax Error | GS | TPR (%) |
| cli | 51 | 0 | 0 | 1 | 0 | 5 | 6 | 0 | 0 |
| codec | 136 | 0 | 1 | 7 | 0 | 28 | 46 | 4 | 0 |
| csv | 33 | 0 | 3 | 0 | 0 | 6 | 8 | 0 | 0 |
| exec | 54 | 0 | 0 | 0 | 0 | 1 | 4 | 0 | 0 |
| fast-pfor | 85 | 1 | 1 | 1 | 0 | 12 | 30 | 1 | 0 |
| fileupload | 43 | 0 | 0 | 0 | 0 | 2 | 4 | 0 | 0 |
| graph | 159 | 0 | 6 | 3 | 0 | 0 | 17 | 3 | 0 |
| jansi | 30 | 0 | 0 | 0 | 0 | 3 | 10 | 0 | 0 |
| pool | 72 | 0 | 1 | 0 | 0 | 4 | 8 | 0 | 0 |
| validator | 119 | 0 | 3 | 0 | 0 | 10 | 19 | 0 | 0 |
| Total | 782 | 1 | 15 | 12 | 0 | 71 | 152 | 8 | 0 |
8.2. Code translation using LLMs.
@@ 9. Threats to Validity
External Validity. One of the key external threats is the generalizability of AlphaTrans. We built and evaluated the first version of AlphaTrans to translate from Java to Python. However, our pipeline is generic, and with minimal effort, the current implementation can translate Java programs to more target languages (e.g., languages supported by GraalVM). Furthermore, the majority of the tools that we used support a large set of programming languages such as JavaScript, Ruby, C/C++, and Rust. Another threat is that AlphaTrans removes certain third-party libraries before translation. Translating repository-level code along with their libraries, especially when the target language does not have equivalent libraries, is a challenging problem that we plan to address in our future work.
Internal Validity. One threat can be the manual validation of the translated types. To address that, several authors have verified the types individually and consulted API documents when necessary. Another threat is that, while all successes reported by the GraalVM validation are true successes, we may have underestimated the capabilities of AlphaTrans by a considerable margin due to the significant proportion of errors caused by limitations in the GraalVM validation approach. To mitigate this threat, we manually augment the universal type map to support a comprehensive translation of types. That said, we have implemented driver code templates to provide a mechanism for adding the support for more types by the users of AlphaTrans if needed.
@@ 10. Concluding Remarks
In this paper, we introduced AlphaTrans. This neuro-symbolic approach combines the power of static analysis and abilities of LLMs in code synthesis to automate repository-level code translation and validation. AlphaTrans decomposes the program into smaller fragments and translates the fragments in the reverse call order, originally building the project in the target language. In addition to syntax check, AlphaTrans implements two levels of validation through GraalVM and test translation. AlphaTrans is the first approach to translate and validate the entire repository, and we envision several research directions to advance repository-level code translation and validation.
One of the major challenges of repository-level code translation is identifying suitable library APIs in the target PL. Often, equivalent Python APIs may not exist, requiring new code generation or translation of the library API itself. Even if similar libraries exist, the logic of libraries might be different in two PLs. AlphaTrans supports translating frequently used APIs and aims to build a generic pipeline. Supporting all the libraries in the pipeline remains an open challenge we aim to address in future work. Furthermore, while the idea of compositional translation and validation is PL-agnostic, the static analysis makes the extension of AlphaTrans to translating from other source projects challenging. Devising LLM-enabled or PL-agnostic static analysis approaches can benefit code translation approaches such as AlphaTrans. We also showed that the quality of the source project test suite can significantly impact the translation validation results. As part of our future work, we plan to integrate an LLM-based test generator into the AlphaTrans pipeline to advance the validation component.